In the laboratory a student finds that it takes 3.19 103 Joules to increase the temperature of 12.s grams of gaseous hydrogen from 20.2 to 38.8 degrees Celsius. The specific heat of hydrogen calculated from her data is JgoC
Answer #1

Q = m * C *delta T

where, Q = heat energy

m = mass

c = specific heat capacity

delta T = T2 - T1

3.19*103 = 12.8 * C * (38.8 - 20.2)

3.19*103 = 12.8 * C * 18.6

3.19*103 = 238.08 * C

C = (3.19*103) / 238 = 0.0134*10^3 J/g0C

Therefore, specific heat = 0.0134*10^3 J/g0C